The intensity of sound from a point source is 1.0 x 10–8 W/m2 at a distance of 5.0 from the source.
What will be the intensity at a distance of 25 m from the source?

Grade:12

1 Answers

Arun
6 years ago
sound intensity  is inversely proportional to r2.
hence
S1/ S2  =  (r2/ r1)2
1 * 10-8 / S2  = (25 / 5)2
S2 = 1* 10 –8/ 25
S2 = 4 * 10 –10  W/m2